14. The total cost function of a firm is given by c = 100 + 40 q + 5q2.



A. Find Total fixed cost



B. Find the Marginal cost equation and the marginal cost of the 5th unit



C. Find Average variable cost equation and the AVC of the 4th unit



D. Find average fixed cost equation and the AVC of the 6th unit



E. Find average total cost equation and the ATC of the 3rd u




1
Expert's answer
2022-02-13T11:50:31-0500

Q=1


a."c=100+40\times1+5(1)^2=145"


b.

"MC=c'=(100+40q+5q^2)'=40+10q"

Q=5

MC=90


c.

"AVC=\frac{(q+5q^2)}{q}=1+5q"

Q=4

AVC=24


d.

"ATC=\frac{100}{q}"

ATC=16.67

AVC=36


e.

Q=3

"c=100+40\times1+5(1)^2=100+40×3+5(3)^2= 265"

ATC=33.33